# Low-energy $N\phi$ scattering from a pole-enhanced triangle diagram

### v1

@@ I Introduction
The second term on the right-hand side of Eq. (19) is proportional to $\vec{P}_{1}^{\,2}$ and vanishes in the limit $\vec{P}_{1}^{\,2}\to 0$. Consequently, only the first term contributes to the leading-order S-wave potential, which reads
$V_{LO}^{s}$ $=$ $-\int_{0}^{\infty}\frac{4\pi\,g_{\rm eff}\,d|\vec{l}|}{8\pi^{3}}\frac{\vec{l}^{\,2}\cdot 8m_{K}}{8\,m_{1}m_{2}m_{3}\left(2m_{\phi}-4\omega_{1}\right)}$ (20)
$=$ $-\frac{1}{4\pi^{2}}\,\frac{m_{K}}{m_{1}m_{2}m_{3}}\int_{0}^{\infty}\frac{g_{\rm eff}\,\vec{l}^{\,2}\,d|\vec{l}|}{m_{\phi}-2\omega_{1}},$
where $\omega_{1}=m_{K}+\vec{l}^{\,2}/(2m_{K})$ has been used.
The momentum integral in Eq. (20) shows a linear ultraviolet (UV) divergence. To regularize this divergence, a hard cutoff $\Lambda$ is introduced, leading to
$\int_{0}^{\Lambda}\frac{\vec{l}^{\,2}\,d|\vec{l}|}{m_{\phi}-2m_{K}-\vec{l}^{\,2}/m_{K}}=\int_{0}^{\Lambda}\frac{\vec{l}^{\,2}\,d|\vec{l}|}{\delta-\vec{l}^{\,2}/m_{K}}$
$=$ $m_{K}\!\int_{0}^{\Lambda}\frac{\vec{l}^{\,2}-m_{K}\delta+m_{K}\delta}{m_{K}\delta-\vec{l}^{\,2}}\,d|\vec{l}|$
$=$ $m_{K}\!\left[-\Lambda-\frac{m_{K}\delta}{\sqrt{-m_{K}\delta}}\arctan\!\left(\frac{\Lambda}{\sqrt{-m_{K}\delta}}\right)\right],$
$=$ $\tilde{I}^{div.}(\Lambda)+\tilde{I}^{con.}(\delta),$
where $\delta=m_{\phi}-2m_{K}$. The first term is linearly divergent in the cutoff $\Lambda$. In contrast, the second term is finite and encodes the non-analytic dependence on the difference of the $K\bar{K}$ threshold to $\phi$ mass to $\phi$ with unphysical masses in the Table. 1. The leading-order potential $V_{LO}^{s}$ is separated into a divergent piece $V^{\rm div}$ and a convergent piece $V^{\rm con}$. The divergent contribution depends on the cutoff $\Lambda$ in the triangle diagram, while the convergent part represents the long-distance contribution from the two-kaon–exchange. Although the numerical value of the $V_{LO}^{s}$ depends on the choice of regularization scheme, the convergent $V^{\rm con}$ depending on the binding energy $\delta$ is regulator-independent.
In addition to employing a momentum cutoff $\Lambda$, the loop integral in Eq. (20) can also be regularized using smooth regulators, such as the monopole or Gaussian form factors. The monopole regulator is expanded as
$\frac{\Lambda_{m}^{2}}{\Lambda_{m}^{2}+\vec{l}^{\,2}}$ $=$ $1-\frac{\vec{l}^{\,2}}{\Lambda_{m}^{2}}+\mathcal{O}\!\left(\frac{\vec{l}^{\,4}}{\Lambda_{m}^{4}}\right),$ (22)
where $\Lambda_{m}$ denotes a hard scale. To leading order in the momentum expansion, this regulator reproduces the behavior of a sharp cutoff up to terms that can be absorbed into higher-order contact interactions.
Alternatively, adopting a Gaussian regulator $\exp(-\vec{l}^{\,2}/\Lambda_{G}^{2})$ leads to a simple mapping between the cutoff scales,
$\Lambda_{G}=\sqrt{\frac{2}{\pi}}\,\Lambda,$ (23)
such that the Gaussian and sharp-cutoff schemes yield equivalent results for the leading contribution to the potential.
Since the study focuses on the long-distance dynamics associated with near-threshold kaons, we do not attempt to resolve the origin of the short-distance contributions. Their effects are encoded in local counter-terms whose values must be fixed by external input and therefore do not affect the qualitative threshold behavior in this work. Therefore, the conclusions regarding the role of triangle diagram and the threshold behavior of the $N\phi$ scattering length are not altered by different regularization prescriptions.
The second term on the right-hand side of Eq. (II) is finite and appears in the form of an $\arctan$ function. Regarding $g_{\rm eff}=g^{2}V^{WT}$, the Weinberg-Tomozawa interaction enters both the divergent and convergent parts of the loop integral. The divergent part is
$V^{\rm div}$ $=$ $-\frac{g^{2}}{4\pi^{2}}\,\frac{m_{K}^{2}}{m_{1}m_{2}m_{3}}\int_{0}^{\Lambda}V^{WT}(|\vec{l}|)\,d|\vec{l}|,$ (24)
which leads to a sum on permutations of $K$ and $\bar{K}$ in the triangle loop, where the Weinberg-Tomozawa interaction corresponds to the isospin channels
$V^{WT}_{I=0,\,N\bar{K}},\quad V^{WT}_{I=1,\,N\bar{K}},\quad V^{WT}_{I=0,\,NK},\quad V^{WT}_{I=1,\,NK},$ (25)
with the corresponding coefficients 3-3, 1-1, 0, and $2$, respectively. The isospin-averaged interactions are given by $V^{WT}_{N\bar{K}}=\frac{1}{2}V^{WT}_{I=0,\,N\bar{K}}+\frac{1}{2}V^{WT}_{I=1,\,N\bar{K}}$ and $V^{WT}_{NK}=\frac{1}{2}V^{WT}_{I=1,\,NK}$.
If dynamically generated poles appear in the $N\bar{K}$ or $NK$ scattering, the Weinberg-Tomozawa interaction should, in principle, be replaced by a resummed series of bubble diagrams, which accounts for multiple rescattering in the corresponding two-body channels Oller et al. (2000). Such a resummation generates near-threshold poles, most notably the $\Lambda(1405)$ in the $N\bar{K}$ scattering, and substantially enhances the triangle amplitude.
To illustrate how these effects enter beyond the Weinberg-Tomozawa vertex, the corresponding two-loop diagrams are shown in Fig. 2. Diagram (a) represents the contribution in which the $N\bar{K}$ (or $NK$) interaction in the triangle is dressed by the resummed bubble series, giving rise to a dynamically generated pole, whereas diagram (b) corresponds to a non-resonant two-loop correction without such pole enhancement.
Once the bubble resummation is inserted into the triangle diagram, both diagrams (a) and (b) generate ultraviolet divergences of $\mathcal{O}(\Lambda^{2})$. These contributions are beyond the renormalization accuracy of the one-loop treatment and must be absorbed into higher-order counter-terms, which is not the focus of this study. Therefore, at order $\mathcal{O}(\Lambda)$, the divergent part $V^{\rm div}$ is consistently evaluated using only the Weinberg-Tomozawa interaction without resummation.
In contrast, the convergent part $V^{\rm con}$ is dominated by low-momentum regions of the loop integration and is sensitive to near-threshold dynamics. In this case, the enhancement induced by the dynamically generated $N\bar{K}$ pole in diagram (a) constitutes a genuine long-distance effect that is retained explicitly, while the contribution from diagram (b) remains subleading and does not involve threshold effect.
As a consequence, the divergent part $V^{\rm div}$ is evaluated with the Weinberg-Tomozawa vertex only, while the effect of the dynamically generated near-threshold pole is incorporated in the convergent part $V^{\rm con}$. Importantly, introducing the pole through the resummed bubble series modifies only the low-energy behavior of the integrand, and does not alter the convergence of the loop integral in Eq. (II). Therefore, it is not necessary to re-evaluate the full triangle-loop integral with dressed propagators in $V^{\rm con}$.
The upper limits of the loop momentum are constrained by the mass at which additional degrees of freedom become relevant. In the case of $N\bar{K}$ rescattering, the $\Lambda(1520)$ resonance is not included explicitly. Accordingly, the loop momentum carried by the $\bar{K}$ in the Weinberg-Tomozawa interaction is restricted by the this excitation, leading to $l^{\rm max}_{N\bar{K}}=245~\mathrm{MeV}$. Similarly, for $NK$ rescattering, the $NK^{\ast}$ channel is not included, and the corresponding loop momentum is bounded by the $NK^{\ast}$ threshold, $l^{\rm max}_{NK}=553~\mathrm{MeV}$.
Focusing on $N\phi$ scattering in the vicinity of the threshold, the momentum dependence of the Weinberg-Tomozawa interaction entering the divergent part $V^{\rm div}$ can be expanded as
$V^{WT}(|\vec{l}|)$ $=$ $V^{WT}(|\vec{l}|=0)+\mathcal{O}(\delta).$ (26)
The $\mathcal{O}(\delta)$ terms are suppressed in the near-threshold region and are neglected at the present order. As a result, the divergent contribution is written as
$V^{\rm div}$ $=$ $-\frac{g^{2}}{4\pi^{2}m_{K}^{2}}\,\tilde{I}^{\rm div}(\Lambda)\,V^{WT}$ (27)
$=$ $\frac{g^{2}}{4\pi^{2}m_{K}}\sum_{i=NK,\,N\bar{K}}V^{WT}_{i}\,\Lambda_{i},$
where the index $i$ labels the $NK$ and $N\bar{K}$ rescattering processes, with $\Lambda_{NK}=l^{\rm max}_{NK}$ and $\Lambda_{N\bar{K}}=l^{\rm max}_{N\bar{K}}$.
Using the relative strengths of the Weinberg-Tomozawa interactions listed in Eq. (25), the divergent part can be expressed as
$V^{\rm div}$ $=$ $\frac{g^{2}V^{WT}_{I=0,\,N\bar{K}}}{4\pi^{2}m_{K}}\left(\frac{1}{3}\,l^{\rm max}_{NK}-\frac{2}{3}\,l^{\rm max}_{N\bar{K}}\right).$ (28)
This expression represents a residual short-distance contribution after partially cancellation among different isospin channels. Numerically, the combination $\frac{1}{3}l^{\rm max}_{NK}-\frac{2}{3}l^{\rm max}_{N\bar{K}}$ is smaller than the characteristic low-energy scale $\sqrt{m_{K}\delta}$ appearing in Eq. (II), indicating that $V^{\rm div}$ remains subleading compared to $V^{\rm con}$.
As a consequence, the triangle diagram leads to
$\frac{1}{V^{\rm div}+V^{\rm con}}$ $\simeq$ $\frac{1}{V^{\rm con}}\left(1-\frac{V^{\rm div}}{V^{\rm con}}\right),$ (29)
where the expansion is justified by $|V^{\rm div}|\ll|V^{\rm con}|$. The subleading term proportional to $V^{\rm div}/V^{\rm con\,2}$ renormalizes the ultraviolet divergence appearing in the $N\phi$ loop function $G$ in Eq. (3). This renormalization is expressed in terms of an effective subtraction scale $\Lambda_{0}=-\pi V^{\rm div}/V^{\rm con\,2}$. According to the parameters adopted in the next section, $\Lambda_{0}$ is found to be about $150~\mathrm{MeV}$ with inputs in Table. 1, which lies well within the momentum range relevant for near-threshold $N\phi$ scattering. This indicates that the effective-range expansion remains valid here, and that corrections beyond the scattering lengths are suppressed. Accordingly, the $N\phi$ scattering lengths in the spin-$1/2$ and spin-$3/2$ channels, which are degenerated at leading order, are approximated as
$-\frac{1}{a_{N\phi}}$ $=$ $\frac{1}{V^{\rm con}}.$ (30)
This relation works as long as the typical external momenta are much smaller than the inverse range of the interaction. The scale is associated with the triangle mechanism, so that the momentum expansion underlying the effective-range expansion is justified in the near-threshold region. Consequently, higher-order terms beyond the scattering length, such as the effective-range parameter, are suppressed.
We emphasize that the present analysis is focused on the threshold behavior of the $N\phi$ amplitude and does not aim at describing possible bound or virtual states close to the threshold, where higher-order terms in the effective-range expansion could become important. Within the kinematic window relevant for the extraction of the scattering length, however, the approximation in Eq. (3) remains reliable.
III Results and discussion
To evaluate the convergent contribution $V^{\rm con}$, the role of near-threshold poles in the $N\bar{K}$ scattering is considered. In the isoscalar channel, the $\Lambda(1405)$ resonance is widely interpreted as a dynamically generated $N\bar{K}$ bound state  Lu et al. (2023); Bulava and others (2024b, a); Guo et al. (2023); Zhuang et al. (2024); He et al. (2024). The proximity of this pole to the $N\bar{K}$ threshold leads to a significant enhancement of the low-energy $N\bar{K}$ amplitude, which in turn promotes the convergent part of the triangle-diagram contribution to $N\phi$ scattering. In the isovector $N\bar{K}$ channel, a virtual pole has been reported in several analyses and is sometimes associated with a broad $\Sigma$ excitation around $1.4~\mathrm{GeV}$ Ma and others (2023); Lu et al. (2023); Lyu et al. (2024); Wang et al. (2024); Li et al. (2024). This pole is located further from the physical sheet than the $\Lambda(1405)$ and therefore provides a subleading correction to the triangle diagram at threshold. By contrast, the $NK$ interaction in the $I=0$ and $I=1$ channels is governed by a vanishing ($C_{ij}=0$) and a repulsive ($C_{ij}=2$) Weinberg-Tomozawa term, respectively, and does not generate a near-threshold pole. As a result, the $NK$ subsystem does not induce an analogous enhancement of the convergent triangle contribution. These considerations indicate that the near-threshold structure associated with the $\Lambda(1405)$ provides the leading enhancement mechanism for $V^{\rm con}$, while other channels yield perturbative corrections.
@@ III Results and discussion
$K$ $524.7$ 495.6
$\phi$ $1048.0$ 1019.5
NN $954.0$ 938.9
@@
$\delta\equiv m_{\phi}-2m_{K}=-\,\tilde{\delta}=-1.0(4)~\mathrm{MeV}.$
@@
The use of an unphysical kaon mass also induces a corresponding modification of the $\phi K\bar{K}$ coupling constant $g$. Invoking $SU(3)$ flavor symmetry, this coupling is related to the $\rho\pi\pi$ coupling, $g_{\rho\pi\pi}$, governing the decay $\rho\to\pi\pi$. The pion-mass dependence of this coupling can be parametrized as
$g_{\rho\pi\pi}=\tilde{c}_{0}+\tilde{c}_{1}\,m_{\pi}^{2}.$
@@
Lattice and phenomenological studies show that $g_{\rho\pi\pi}$ varies only mildly, at the level of $2\%\!-\!4\%$, when the pion mass is varied from $140$ to $200~\mathrm{MeV}$ Wang et al. (2025), in agreement with earlier analyses indicating an approximate pion-mass independence of this coupling Hanhart et al. (2008); Guo et al. (2011). Consequently, the modification of the $\phi K\bar{K}$ coupling induced by the unphysical light-quark masses is expected to be numerically small and does not affect the qualitative features of the near-threshold $N\phi$ interaction discussed below.
Subsequently, the leading-order S-wave potential induced by the triangle diagram is estimated analytically as
[…]
which is a scalar quantity and provides the large contribution to the $N\phi$ scattering length. The factor $R_{\rm pole}$ parametrizes the relative enhancement of the triangle amplitude induced by the near-threshold $\Lambda(1405)$ pole compared to the leading-order Weinberg-Tomozawa interaction,
$R_{\rm pole}$ $=$ $\frac{g_{i}^{2}/\bigl|\,th_{N\bar{K}}-\sqrt{s_{\rm pole}}\,\bigr|}{V^{WT}}=5.65^{+1.13}_{-0.96},$ (32)
with the $N\bar{K}$ threshold $th_{N\bar{K}}$, where $\sqrt{s_{\rm pole}}$ denotes the pole position of the $\Lambda(1405)$, and the coupling $g_{i}$ is taken from Ref. Guo and Oller (2013). The quoted uncertainty reflects the difference between the Fit-I and Fit-II parameterizations.
It is worth emphasizing that the pole contribution is qualitatively different from a local contact interaction such as $V^{WT}$. Owing to the proximity of the $\Lambda(1405)$ pole to the $N\bar{K}$ threshold and its strong coupling to the $N\bar{K}$ channel, the corresponding amplitude shows a pronounced energy dependence as the center-of-mass energy approaches $\sqrt{s_{\rm pole}}$. Although $s_{\rm pole}$ is complex, the $\Lambda(1405)$ remains relatively narrow and dominantly couples to the $N\bar{K}$ channel, while its couplings to other channels are not dominant Lu et al. (2023); Bulava and others (2024b, a); Guo et al. (2023); Zhuang et al. (2024); He et al. (2024).
In the low-energy $N\phi$ scattering, the kinematics constrain the intermediate $N\bar{K}$ subsystem to energies close to the $N\bar{K}$ threshold. The enhancement induced by the $\Lambda(1405)$ pole is captured by the factor $R_{\rm pole}$, which governs the main contribution of the $N\bar{K}$ rescattering to the triangle diagram.
@@
The numerical results for the $N\phi$ scattering length are obtained by varying the binding energy of the intermediate $K\bar{K}$ system to the $\phi$ meson. The corresponding results are shown in Fig. 3. The calculated curves show good agreement with the magenta band extracted in the spin-$3/2$ channel at an unphysical pion mass in Ref. Lyu et al. (2022), as well as with the green band corresponding to the spin-averaged scattering length measured by the ALICE collaboration Acharya and others (2021).
@@
The uncertainty associated with the enhancement factor $R_{\rm pole}$, which is determined from the physical parameters and pole couplings, propagates into the prediction of the scattering length. This effect is illustrated in Fig. 4, where the results are presented for the central value of $\tilde{\delta}$, and the uncertainty bands reflect the variation of $R_{\rm pole}$. The theoretical curves are fully covered by the corresponding uncertainty bands, indicating a stable prediction within the adopted parameter ranges.
@@
Figure 3: The $N\phi$ scattering length from two-Kaon-Exchange with $\tilde{\delta}=\left(1.0\pm 0.2\right)\,\rm{MeV}$, where the solid, dashed, and dotted lines correspond to the central value, lower and upper limits of $\delta$, respectively. The magenta and green bands represent the real part of the scattering lengths from HAL QCD Lyu et al. (2022) and ALICE collaboration Acharya and others (2021), respectively.
@@
Figure 4: The $N\phi$ scattering length from two-Kaon-Exchange with uncertainty in $R_{pole}$ deduced by physical couplings.
@@
Concerning the use of physical couplings in $V^{\rm con.}$, the possible modification of the coupling $g_{i}$ induced by the unphysical hadron masses is estimated at the order-of-magnitude level by a mass extrapolation,
@@
which serves only as an indicative estimate. A quantitative determination of this dependence would require dedicated lattice QCD input and is beyond the scope of this study.
The uncertainty in $g_{i}$ corresponds to an overall variation of the potential at the level of $\pm 19\%$. As a consequence, the uncertainty bands shown in Fig. 5 are comparable to those in Fig. 4, with the dashed and dotted curves remaining within the corresponding error bands. This stability indicates that the triangle-diagram–driven dynamics persists once the quark-mass dependence of the couplings is considered. Therefore, the conclusion on the dominance of the two-kaon–exchange triangle mechanism remains unaffected by reasonable variations of the coupling constants.
Figure 5: The $N\phi$ scattering length from two-Kaon-Exchange with uncertainty in $g_{i}$.
@@
where $n_{0}=2$ and $4$ correspond to the cases in which the coupling $g$ is taken to be independent of, or dependent of, $F_{\pi}$, respectively.
The variations of the potential translate into the scattering lengths shown in Fig. 6, where the dashed and dotted curves indicate the corresponding upper and lower limits. These curves remain within the cutoff-variation bands for $\Lambda\in[20,300]\,\mathrm{MeV}$, demonstrating a good numerical convergence. The narrow separation between the dashed and dotted curves further confirms that the pion-mass dependence of $F_{\pi}$ is perturbative. Therefore, the uncertainty associated with the light-quark-mass dependence of $F_{\pi}$ does not affect the conclusions.
Figure 6: The $N\phi$ scattering length from two-Kaon-Exchange with uncertainty in $F_{\pi}$, where the solid, dashed, and dotted lines correspond to the central value, lower, and upper limits of $n_{0}=0,\,2,\,4$, respectively.
Figure 7: The $N\phi$ scattering length from Two-Kaon-Exchange with choices on $\tilde{\delta}$, where the solid, dashed and dotted lines correspond to the central value, lower and upper limits of $\delta$, respectively.
[…]
Beyond the overall agreement of the scattering length with lattice-QCD and femtoscopy results, a particularly distinctive feature of the present analysis is the threshold behavior of the leading-order potential $V_{LO}^{s}$. Near the $N\phi$ threshold, $V_{LO}^{s}$ shows a power-law dependence on the the difference of the $K\bar{K}$ threshold to $\phi$ mass $\tilde{\delta}$, with $V_{LO}^{s}\propto\tilde{\delta}^{\,n}$. The exponent interpolates between $n=0$ for small cutoff values and $n=1/2$ for large cutoffs, reflecting the interplay between the near-threshold kinematics and the loop momentum region probed by the triangle diagram. This behavior is different from that expected for long-range interactions. In particular, the QCD van der Waals force generated by multi-gluon exchange leads to a much steeper dependence, $V\propto\delta^{3}$ or $\delta^{7/2}$ (equivalently $p_{N\phi}^{6}$ or $p_{N\phi}^{7}$) Appelquist and Fischler (1978); Brambilla et al. (2017). The power-law scaling found here also differs from that associated with one-boson-exchange descriptions or Yukawa-type interactions Kim et al. (2021), and can be further tested in the three-body system Wen et al. (2025).
In the case $n=0$, which corresponds to a small cutoff $\Lambda$, the $N\phi$ scattering length becomes only weakly dependent on the difference of the $K\bar{K}$ threshold to $\phi$ mass. This behavior indicates that the system approaches a regime dominated by near-unitary three-body $NK\bar{K}$ dynamics, in which the characteristic length scale is set by the large scattering length rather than by the range of the underlying interaction. Such a regime provides a particularly sensitive window to three-body correlations, which may act as an ultra-long-range component of the effective interaction.
In this extreme low-energy limit, subleading effects-such as electromagnetic interactions among hadrons-may become comparatively more relevant than in typical hadronic systems, as discussed in Refs. Gasser et al. (2008); Baru et al. (2011a, b); Zhang and Guo (2021); Shi et al. (2022). Their quantitative impact, however, depends on the proximity to the unitary regime and requires a dedicated analysis beyond the scope of this study.
For $n=1/2$, realized for larger cutoff values, the $N\phi$ scattering length shows an approximately linear dependence on the mass difference between the $\phi$ meson and the $K\bar{K}$ threshold. Both regimes highlight the sensitivity of the $N\phi$ interaction to threshold kinematics and three-body dynamics. A systematic exploration of these behaviors calls for high-precision scattering data and is well-suited for future lattice-QCD simulations with variations of the light-quark masses. In this sense, the $N\phi$ system offers a rare example in which threshold three-body dynamics leaves a direct imprint on a two-body observable.
@@
On the other hand, for physical hadron masses one finds $\sqrt{m_{K}|\tilde{\delta}|}/m_{K}\simeq 0.24$, indicating that the kaons propagating in the triangle diagram remain non-relativistic. This provides a self-consistency check for the expansion in Eq. (31) and motivates an evaluation of the $N\phi$ scattering length using physical masses. The real part of the scattering length is shown in Fig. 8, where the momentum carried by the outgoing Kaon is $|\vec{k}_{K}|=118.17\,\mathrm{MeV}$.

### v2

@@ I Introduction
The second term on the right-hand side of Eq. (10) is proportional to $\vec{P}_{1}^{\,2}$ and vanishes in the limit $\vec{P}_{1}^{\,2}\to 0$. Consequently, only the first term contributes to the leading-order S-wave potential, which is convergent and reads
$V_{LO}^{s}$ $=$ $\int\frac{d^{3}l}{(2\pi)^{3}}\,\frac{g_{\rm eff}}{8\,m_{K}^{3}}\,\frac{4\vec{l}^{2}}{(P_{1}^{0}-\omega_{1}-\omega_{2})^{2}}$ (11)
$\times\frac{u(P_{2})\,\bar{u}(P_{4})}{2m_{\Lambda}\left(P_{1}^{0}+P_{2}^{0}-\omega_{1}-\omega_{\Lambda}\right)},$
$=$ $\int\frac{d^{3}l}{(2\pi)^{3}}\,\frac{g_{\rm eff}}{8\,m_{K}^{3}}\,\frac{4\vec{l}^{2}u(P_{2})\,\bar{u}(P_{4})}{\left(\delta-\frac{\vec{l}^{2}}{m_{K}}\right)^{2}}\frac{1/2m_{\Lambda}}{\Delta E-\frac{\vec{l}^{2}}{2m_{\Lambda}}}$
where $\delta=m_{\phi}-2m_{K}$ and $\Delta E=m_{\phi}+m_{p}-\omega_{1}-m_{\Lambda}$ with unphysical masses in the Table. 1.
Unphysical case: With the definitions $a_{0}=\sqrt{-m_{K}\delta}$ and $b^{2}=-2m_{\Lambda}\Delta E$, the momentum integral in Eq. (11) is evaluated using a parameter $\Lambda$, yielding
$\frac{V^{s}_{LO}}{\mathcal{C}}$ $=$ $\arctan\left(\frac{\Lambda}{a_{0}}\right)\left[\frac{a_{0}}{2(b^{2}-a_{0}^{2})}+\frac{a_{0}(a_{0}^{2}-2b^{2})}{(a_{0}^{2}-b^{2})^{2}}\right]$ (12)
$+\frac{a_{0}}{2(b^{2}-a_{0}^{2})}\frac{a_{0}\Lambda}{a_{0}^{2}+\Lambda^{2}}$
$+\arctan\left(\frac{\Lambda}{b}\right)\frac{b^{3}}{(b^{2}-a_{0}^{2})^{2}},$
where $\mathcal{C}=-2m_{K}^{2}g_{\rm eff}/\pi^{2}$. It is worth noting that the momentum integral in Eq. (11) is strictly UV convergent by simple power counting. Therefore, the parameter $\Lambda$ introduced here is not a mathematical regularizer for a divergence, but rather a physical scale characterizing the finite size of the interacting hadrons. The analytical expressions reveal that the potential $V_{LO}^{s}$ is significantly enhanced as $|a_{0}|$ and $|b|$ vanish, corresponding to the $K\bar{K}$ threshold approaching the $\phi$ mass and the $\Lambda(1405)$ state becoming on-shell. This behavior is a signature of the kinematical enhancement inherent in the pole-promoted triangle-like mechanism.
In the limit where $|b|$ acts as a hard scale relative to $|a_{0}|$, the interaction simplifies to
$\frac{V^{s}_{LO}}{\mathcal{C}}$ $\simeq$ $\frac{-3a_{0}}{2b^{2}}\arctan\left(\frac{\Lambda}{a_{0}}\right)+\frac{1}{b}\arctan\left(\frac{\Lambda}{b}\right)$ (13)
$+\frac{a_{0}}{2b^{2}}\frac{a_{0}\Lambda}{a_{0}^{2}+\Lambda^{2}},$
which is proportional to $a$ within a narrow energy window. For the hierarchy $\Lambda\gg|b|\gg|a|$, the interaction reads
$\frac{V^{s}_{LO}}{\mathcal{C}}$ $\simeq$ $\frac{\pi}{2}\left(\frac{2b-3a_{0}}{2b^{2}}\right),$ (14)
implying that the $\Lambda(1405)$ plays a crucial role in $N\phi$ scattering. This behavior differs from that characteristic of long-range interactions. In particular, the QCD van der Waals force arising from multi-gluon exchange exhibits a much steeper dependence, $V\propto p_{N\phi}^{6}$ or $p_{N\phi}^{7}$ Appelquist and Fischler (1978); Brambilla et al. (2017). The derived interaction is also distinct from one-boson-exchange models or Yukawa-type interactions Kim et al. (2021) and can be further tested in three-body systems Wen et al. (2025).
@@
$K$ $524.7(2)$ 495.6
$\phi$ $1048.0(4)$ 1019.5
NN $954.0(2.9)$ 938.9
Physical case: For physical hadron masses, the ratio $\sqrt{m_{K}|\tilde{\delta}|}/m_{K}\simeq 0.24$ indicates that the intermediate Kaons in the triangle diagram remain non-relativistic. This validates the expansion in Eq. (11) and motivates the evaluation of the $N\phi$ scattering length using physical parameters. In this regime, $\delta$ becomes positive, introducing a pole in the propagator at momentum $k_{0}=\sqrt{2m_{K}\delta}$. Consequently, the interaction takes the form
$\frac{V_{LO}^{s}}{\mathcal{C}}$ $=$ $\frac{k_{0}}{4\left(b^{2}+k_{0}^{2}\right)}\left(\log\frac{\Lambda+k_{0}}{\Lambda-k_{0}}-i\pi-\frac{2k_{0}\Lambda}{\Lambda^{2}-k_{0}^{2}}\right)$ (15)
$+\frac{k_{0}\left(k_{0}^{2}+2b^{2}\right)}{2\left(b^{2}+k_{0}^{2}\right)^{2}}\left(\log\frac{\Lambda-k_{0}}{\Lambda+k_{0}}+i\pi\right)$
$+\frac{b^{3}}{\left(b^{2}+k_{0}^{2}\right)^{2}}\arctan\left(\frac{\Lambda}{b}\right),$
where, in the large-$\Lambda$ limit, the real part scales as $1/b$. This behavior aligns with the result for unphysical hadron masses in Eq. (12), confirming that the physical $N\phi$ scattering driven by the triangle mechanism is distinct from the QCD van der Waals force. The imaginary part, arising from the $\pm i\pi$ terms, reflects the open-channel effect of the $\phi\to K\bar{K}$ decay characterized by $k_{0}$. Furthermore, the $\Lambda$ dependence is logarithmic and remains mild for $\Lambda\gg k_{0}$.
III Results and discussion
@@
$\delta\equiv m_{\phi}-2m_{K}=-\,\tilde{\delta}=-1.4(8)~\mathrm{MeV}.$
@@
The use of an unphysical Kaon mass also induces a corresponding modification of the $\phi K\bar{K}$ coupling constant $g$. Invoking $SU(3)$ flavor symmetry, this coupling is related to the $\rho\pi\pi$ coupling, $g_{\rho\pi\pi}$, governing the decay $\rho\to\pi\pi$. The pion-mass dependence of this coupling can be parametrized as $g_{\rho\pi\pi}=\tilde{c}_{0}+\tilde{c}_{1}\,m_{\pi}^{2}.$ Lattice and phenomenological studies show that $g_{\rho\pi\pi}$ varies only mildly, at the level of $2\%\!-\!4\%$, when the pion mass is varied from $140$ to $200~\mathrm{MeV}$ Wang et al. (2025), in agreement with earlier analyses indicating an approximate pion-mass independence of this coupling Hanhart et al. (2008); Guo et al. (2011). Consequently, the modification of the $\phi K\bar{K}$ coupling induced by the unphysical light-quark masses is expected to be numerically small and does not affect the qualitative features of the near-threshold $N\phi$ interaction discussed below.
@@
The numerical results for the $N\phi$ scattering length are obtained by varying the binding energy of the intermediate $K\bar{K}$ system to the $\phi$ meson, where the center values of $\sqrt{s}_{R}$ and $g_{i}$  Guo and Oller (2013) are adopted. The corresponding results are shown in Fig. 4, where the dash-dot-dot line represents the case of switching off the imaginary part of $\sqrt{s}_{R}$. It is easy to see that the difference between the dashed and dotted lines is small, which implies that the changes in $\delta$ affect the interaction in Eq. (12) only mildly. On the other hand, the obvious difference between the solid and dash-dot-dot lines is caused by keeping the imaginary part of $\sqrt{s}_{R}$ or not, which supports that $\Lambda(1405)$ is crucial to yield the interaction in the triangle diagram. Additionally, the calculated curves, with $\Lambda>700\,\rm{MeV}$, show good agreement with the magenta band extracted in the spin-$3/2$ channel at an unphysical pion mass in Ref. Lyu et al. (2022), as well as with the green band corresponding to the spin-averaged scattering length measured by the ALICE collaboration Acharya and others (2021).
Subsequently, the leading-order S-wave potential induced by the triangle diagram is estimated analytically in Eq. (12), which is a scalar quantity and provides a large contribution to the $N\phi$ scattering length. The enhancement of the triangle amplitude induced by the near-threshold $\Lambda(1405)$ pole is determined by the pole position $\sqrt{s}_{R}$ and the coupling $g_{i}$ in Ref. Guo and Oller (2013).
@@
The uncertainty associated with $\sqrt{s}_{R}$ and $g_{i}$, which is determined from the physical parameters and pole couplings, propagates into the prediction of the scattering length. This effect is illustrated in Fig. 5, where the results are presented for the central value of $\tilde{\delta}$, and the lines reflect the uncertainty of the $\Lambda(1405)$ pole. The discrepancy between the theoretical predictions is very small and is larger than the cases in Fig. 4, which indicates that the $\Lambda(1405)$ dominates over the $\phi$ coupling to $K\bar{K}$ in the triangle diagram. In the meantime, the uncertainty bands fully cover the theoretical curves.
@@
Figure 4: The $N\phi$ scattering length from two-Kaon-Exchange with $\tilde{\delta}=\left(1.4\pm 0.8\right)\,\rm{MeV}$, where the solid, dashed, and dotted lines correspond to the central value, lower and upper limits of $\delta$, respectively. The magenta and green bands represent the real part of the scattering lengths from HAL QCD Lyu et al. (2022) and ALICE collaboration Acharya and others (2021), respectively.
@@
Figure 5: The $N\phi$ scattering length from two-Kaon-Exchange with uncertainty in $\Lambda(1405)$ pole.
@@
Concerning the use of physical couplings in $V^{s}_{LO}$, the possible modification of the coupling $g_{i}$ induced by the unphysical hadron masses is estimated at the order-of-magnitude level by a mass extrapolation,
@@
which serves only as an indicative estimate. A quantitative determination of this dependence would require a dedicated lattice QCD input is beyond the scope of this study.
The uncertainty in $g_{i}$ corresponds to an overall variation of the potential at the level of $\pm 19\%$. As a consequence, the uncertainties shown in Fig. 6 are comparable to those in Fig. 5, with the dashed and dotted curves remaining within the corresponding error bands. This stability indicates that the triangle-diagram–driven dynamics persists once the quark-mass dependence of the couplings is considered. Therefore, the conclusion on the dominance of the two-Kaon–exchange triangle mechanism remains unaffected by reasonable variations of the coupling constants.
Figure 6: The $N\phi$ scattering length from two-Kaon-Exchange with uncertainty in $g_{i}$ with unphysical hadron masses.
@@
where $n_{0}=2$ and $4$ correspond to the cases in which the coupling $g$ is taken to be independent of, or dependent on, $F_{\pi}$, respectively. This uncertainty is much smaller than the one in Eq. (16) and counted perturbatively.
@@
On the other hand, for physical hadron masses, the Kaons propagating in the triangle diagram remain non-relativistic. This provides a self-consistency check for the expansion in Eq. (11) and motivates an evaluation of the $N\phi$ scattering length using physical masses in Eq. (15). The real part of the scattering length is shown in Fig. 7, where the momentum carried by the outgoing Kaon is $|\vec{k}_{K}|=118.17\,\mathrm{MeV}$.